how to solve a system by elimination?
example : 3x - y = 13 and 3x + y = 17

Found 2 solutions by ewatrrr, sudhanshu_kmr:
Answer by ewatrrr(24785)

 
Hi
solving by elimination
3x - y = 13
3x + y = 17
3x - y = 13
-3x - y = -17 |subtracting 2nd EQ from 1st EQ to eliminate x
-2y = -4
y = 2 and x = 5 (3x-2=13, 3x = 15, x = 5)
CHECKING our Answer***
3x + y = 17
15 + 2 = 17

Answer by sudhanshu_kmr(1152)


add both equations,
you will find 6x = 30 ,
or, x = 30/6 = 5
now, put the value of x on first equation,
3*5 -y = 13
or -y = 13 -15
or y = 2
so, x=5, y= 2